Sol Summary – March 8th

Crew 206 March 8th 2019
Sol 12

Summary Title :

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Since the begining of the week we are waking up
one or two minutes later every day, tomorrow we will have to get back in
track ! We still do our workout sessions as seriously as the first day
though, our scores have greatly increased. With the lack of an EVA we
decided to clean the Hab again, the dust settles everywhere so quickly
that it’s a necessity. We also had some long discussions on the methods
we use to measure the water consuption of flushes, we have to continue
averaging the water volume over many toilet flushes. For lunch we
prepared some pizzas, we ate a wonderfull lunch that made us all happy.
After that we played some Tarot, a popular french card game. In the
afternoon there was a TELEOP session, it worked fine, this experiment is
back on track ! As planned, there was also a VR session, we are
continuing to gather data.

Look ahead plan : We hope to go on EVA tomorrow but the weather doesn’t
seem promising, so we might just continue our indoor experiments in the
next days.

Anomalies in work : None

Weather : Cloudy all day with a few sunshines

Crew physical status : Fine

EVA : None

Reports to be filed : GreenHab, Journalist, Operation, EVA Request is
weather allows it

Support Requested : None

Journalist Report – March 7th

Crew 206 – 03/07/2019

Sol 11

Author: Benjamin Auzou, Journalist

“An unflappable crew”

This mission in the MDRS is far from a long calm river, outside the Hab with the weather that changes everyday and also inside the Hab with the disagreements that can happen or with the experiments issues. However we are always acting like a crew and manage to tackle all situations. As a matter of fact, we are working together for more than one year and we are able to evolve together, understand each other’s point of vue or limits. And above all the mission is working so well because we have a good communication and ability to be serious, but more importantly fun and humour have an essential role in our relationships. Indeed life under confinement is easier, and cohesion is improved when the crew can put things in order with laugh.

After the perturbed comms window of yesterday, we played cards. Jérémy, Aurélien, Norbert and I played Monopoly Deal, our new favorite game of the mission, while Gaspard and Cerise were in a furious game of “Crapette rapide”. Gaspard won all the games, he is as fast at crapette as he is lucky at Monopoly.

For Cerise’s birthday, we had a king’s breakfast, lunch and diner : pancakes, cheddar biscuits with garlic and herbs butter, chicken curry, chocolate cake, mixed salad, pastas with fresh pesto. The rest of the crew prepared a wonderful present for Cerise. With that colorful clutch bag she will be a real fashion trendsetter when she will get back to Paris..

We walked out on EVA to rescue our fallen experiment, the LOAC and the weather station. Arriving next to it with the rovers, we evaluated the damages : a strong wind lifted one foot of the tripod from the ground. After placing the station safe back in the engineering airlock we started our complete check of the MDRS to prepare it for the strong winds to come. We fixed all the tarps of the tunnels. Tomorrow and Saturday, the area will be hit by gale force winds. We are ready for that.

Operations Report – March 7th

Crew 206 Operations Report 07 – MAR – 2019

SOL: 11

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 10 hours

Solar— SOC % – Turned off at 76%, Turned on at 80% (19h00)

SOC tracking :
7h30: 85%
9h30: 92%
11h30: 95%
13h30: 92%
15h30: 88%
16h30: 86%
18h30: 82%

Diesel Reading – 65%
Propane Reading – 78%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 400 gallons

Water differential (static tank) – 13 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 237 gallons

Static to Loft Pump used – Yes

Water Meter: 0141689,7 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Used

Hours: 77,9 hours

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Spirit rover used: Used

Hours: 78,9 hours

Beginning charge: 100%

Ending charge: 96%

Currently charging: Yes

Notes on rovers: Opportunity’s breaking system seems to be weak or broken. More investigation tomorrow with the help of Atila.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Still investigating. Nothing more to report today.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Secured the Hab for tomorrow’s
strong winds.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Astronomy Report – March 7th

Astronomy Report, Sol 11

Name: Aurélien Mure

Crew: 206
Date: 03/07/2019


Features observed : Sun spot

Problem encountered : Little sun spot observed but only 15 minutes of sun, so not enough time to take pictures.


Object observed : None

Problems encountered : I have been waiting for two days for my images of M45 to be taken by the telescope. I’ll my objectives tomorrow. [end]

GreenHab Report – March 7th

Crew 206 Green Hab Report – 07-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (7 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 16.1°C

High: 40.8°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.3 gallons

Water in Blue Tank: 236.8 gallons

Crops watered at 0915 hours (2.15 gallons) and 1830 hours (2.15 gallons):


1) 08.30 AM 25.3° C / 26 %
2) 11.30 AM 24.4° C / 29 %
3) 15.00 PM 20.4° C / 31 %
4) 18.30 PM 21.3° C / 28 %

Changes to crops: None.

– Cress planted this week sprouted only in one of the two gardener, because in one gardener the water couldn’t flow. I resolved it. – Since we harvested some salads for lunch (savory mix), a gardener is empty, so I hoed it and prepared the soil, to plant new lettuces tomorrow.
– The door was opened from 0830 to 1500, and the fan was turned on for almost 3 hours (by short periods of time), but even with that the heat increased in the greenhab to reach 40°.

Harvest: 64 g of savory mix, 22.7 g of radish, 2.7 g of parsley, 1 g of cress, 34 g of basil.

Support/supplies needed: None.

EVA Report – March 7th

Crew 206 EVA Report 07 Mar 2019
EVA #09

Author : Aurélien Mure

Purpose of EVA:
We realized that the weather station and the LOAC had fallen down, so we need an EVA first to assess the damages and to fix it then. We will go out with some tools to be prepared to every case of damage. We won’t have plan B for tomorrow as we know the weather is not fine. Our equipment is endangered so we decide to ask to out.

Start time: 9:15 am

End time: 11:15 am

Narrative: We left the airlock at 9:18 and went to drive Curiosity for team Red (Gaspard and Norbert) and Opportunity for team Blue (Benjamin and me). The brakes of Opportunity were really lazy this morning so team Blue decided to go with Spirit under the HabCom approval. As soon as we arrived at the LOAC location we realized that the weather station has fallen down on its fragile parts. The weather cock and the anemometer are down and appear to be damaged. We quickly decide to take all the equipment back to the Hab to work on it. As strong wind are announced for saturday it is clear that we will deploy the equipment on monday.

Destination: LOAC and Weather Station location, Hab Ridge

Coordinates: 518500/425700 (LOAC and Weather Station), 517900/4251200 (Hab Ridge Road /Sagan Street)

EVA Participants: Norbert Pouzin (GHO), Aurélien Mure (XO), Jeremy Auclair (CMD, EVA leader)

Roads and routs per MDRS Map: Cow Dung Road

Mode of travel: Driving and Walking

Sol Summary – March 7th

Crew 206 March 7th 2019
Sol 11

Summary Title : Preparing for the Gale force winds

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Waking up is becoming harder and harder for me, even if I sleep as much as before. Some of us sleep better and some of us sleep worse than usual, confinement and isolation have different effects on different people. We are nonetheless sticking to our strict daily routine : wake up at 07:00 am, sport and breakfast. The EVA’s purpose was to asses the damage on the weather station (that tipped over yesterday due to the winds) and the LOAC, and to dismantle the instruments to bring them back to the Hab. The weather cock on the station broke, but the sensor is still operationnal. The rest of the equipment is intact. For the rest of the EVA the team attached more tightly all the tarps that cover the tunnels.
Today is Cerise’s birthday, so the crew prepared a very nice lunch, and Cerise prepared a great chocolate cake. We had a very nice time. In the afternoon we were able to run a digital TELEOP session with the new computer station we received, so we are now collecting twice as much data, things all finaly looking good for TELEOP. The wind is sporadically blowing hard on the station, we hope that it won’t be too bad in the next few days.

Look ahead plan : Almost all our experiments are running fine, except the weather station. We will take advantage of the fact that the LOAC is inside to measure the quality of the air we breathe.

Anomalies in work : None

Weather : Sunny between 11 am and 12 am, Cloudy for the rest of the day

Crew physical status : Fine

EVA : LOAC and weather station dismantling

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA Request

Support Requested : Cooking whip and pressurized air can

GreenHab Report – March 6th

Crew 206 Green Hab Report – 06-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (1 hrs)

Shade Cloth (40%)

Average temperatures:
Low: 17.3°C
High: 36.9°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.1 gallons

Water in Blue Tank: 241.1 gallons

Crops watered at 0915 hours (1.3 gallons) and 1830 hours (2.8 gallons):


1) 09.15 AM 18.9° C / 42 %
2) 11.00 AM 22.6° C / 40 %
3) 15.00 PM 27.9° C / 23 %
4) 18.30 PM 21.3° C / 37 %

Changes to crops: Cress keeps sprouting but only on one gardener.

Narrative: Today the greenhab planning was quite disturbed due to the journalists : I did my experiment the morning, but on the afternoon I couldn’t go there to monitor the temperature.

Harvest: Parsley : 3g

Support/supplies needed: None.

Journalist Report – Mar 06th

Crew 206 – 03/06/2019

Sol 10

Author: Benjamin Auzou, Journalist

“Into the Wind”

I spent a very bad night, measured by my Dreem headband as 5 hours 49 minutes night. This device is a sleeping band we wear each night and that monitor our sleep quality : duration, sleep phases, agitation… Confinement has a sneaky infuence on humans. It can change behaviour, impulsivity, irritability or motivation. Personnaly, it seems to have an impact on my sleep !

Sadly we have no pictures of the EVA of the Day, indeed we left the Hab with no SD card in the camera. We have a lot of checks before leaving the airlock : working suit, gloves, med kit, GPS, watch, radios. Unfortunately this morning we were attentive on the security check and forgot that little card ! The walk was beautiful between Canyons, the Reservoir Dam, flat and rocky areas. We shot a lot of pictures with our eyes, sealed in our memories : “Click” as our astronomer and executive officer like to say in such moments


The MDRS Campus has been visited today by a film crew. They were very respectful of our mission and we managed to continue our martian work and life without seeing them any time. Except Jérémy that have been interviewed for the documentary.

In the afternoon, the Hab was hit by the wind. Thanks to our weather station, we were supposed to receive the values of the wind : direction and speed. Unfortunately for us, the wind was too strong for the installation and the tripod fell on the ground. The martian soil is not ideal to fix things with pegs. We will go for a reinstallation and reparation (if necessary) EVA tomorrow. It’s sad when we think that less than three hours before we were next to it changing the batteries. The adaptation to environment is important for missions to Mars. On Earth, the more similar example is science lead in Sahara desert or in Antarctica.

The end of the afternoon was more calm and relaxed. The wind stopped making noises around the MDRS. We were all together in the Hab talking, working on our reports, and watching VR videos for a human factors experiment.

We faced snow at the beginning of the mission, then rain, and now wind is threatening us. The martian climate isn’t clement with us ! But we managed to adapt our mission and our goal each time the weather changed !

Copyright © The Mars Society. All rights reserved. | Main Site